# Sectorial covers over fanifolds

### v2

@@ 4.3. Construction
Suppose that $\Phi$ is not closed and, for instance, there is a half-open $1$-stratum $I$ adjacent to $P$. Recall that in this case the handle attachment $\mathbf{W}(\Phi_{1})\#_{\mathcal{L}_{1}}(T^{*}\widehat{M}_{I}\times T^{*}I_{\circ})$ is not precisely a Liouville manifold-with-boundary, as one needs to modify the canonical Liouville form of $T^{*}I_{\circ}$ near $\partial T^{*}I_{\circ}$ making it convex. This modification, which is not necessary if $I$ were closed, yields a stop on the newly formed ideal boundary of the handle attachment.
@@
We adapt the idea in [GPS2, Example 1.34]. Again, assume that $\Phi$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata. Finiteness of the cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ follows from the condition $(i)$. By construction each $\widetilde{\mathbf{W}}(P)$ has no corners and $\partial\widetilde{\mathbf{W}}(\Phi)$ is empty. Moreover, $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ intersect transversely for any distinct pair $P,P^{\prime}$ of 00-strata unless they are disjoint. By Lemma 4.7 each $\widetilde{\mathbf{W}}(P)$ is a Weinstein sector. Hence the collection of the hypersurfaces $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial if there exist functions
@@
Consider the functions $y_{P,I}$ in the proof of Lemma 4.7. Pushing forward $X_{y_{P,I}}$ by $\bar{\pi}$, we obtain a collection of vector fields on neighborhoods of $\bar{\pi}(H_{P,I})$. Here, by a vector field on a subset of $\Phi$ we mean the gluing of vector fields on submanifolds of $\Phi_{k}\setminus\Phi_{k-1}$ for $k=1,\ldots,n$. Take their deformations around the corners to concatenate $\bar{\pi}_{*}(X_{P,I})$ so that translation of the result gives a vector field $\tilde{X}_{P}$ on a neighborhood $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$, which is outward pointing along $\partial^{2}\widetilde{\mathbf{W}}(P)$. Recall that $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P)),\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P^{\prime}))$ intersect transversely for any distinct pair $P,P^{\prime}$ of 00-strata unless they are disjoint. Since $\Phi$ has only finitely many 00-strata, we may deform further $\tilde{X}_{P}$ simultaneously so that they in addition become tangent to $\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ on the intersections. We use the same symbol $\tilde{X}_{P}$ to denote these deformations. We define $\tilde{y}_{P}$ as the Hamiltonian lifts of $\tilde{X}_{P}$, which make sense due to the construction of $\widetilde{\mathbf{W}}(\Phi)$ and our definition of a vector field on a subset of $\Phi$.
@@
The construction does not work without modification of $\widetilde{\mathbf{W}}(\Phi)$, as in general $\partial\widetilde{\mathbf{W}}(\Phi)$ is nonempty and intersects $\partial^{2}\widetilde{\mathbf{W}}(P)$.
@@ 4.4. Sectorial descent and HMS over fanifolds
which is obtained from $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)}\colon\operatorname{Exit}(\Phi)\to^{*}\operatorname{DG}^{*}$ in [GS23, Section 4.5] by taking adjoints of the images of the morphisms. By definition both $\operatorname{fsh}_{*}$ and $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)*}$ are constructible cosheaves of categories over $\Phi$.
Lemma 4.12 (cf. [GS23, Proposition 4.34]).
There is an isomorphism
@@
Proof.
The claim immediately follows from [GS23, Proposition 4.34]. ∎
Third, let $\operatorname{UI}(\Phi)$ be the category of unions of intersections of $\widetilde{\mathbf{W}}(P)$ in the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ from Lemma 4.10. Morphisms are given by inclusions of Weinstein sectors. Introduce to $\operatorname{UI}(\Phi)$ a topology generated by maximal proper Weinstein subsectors $\widetilde{\mathbf{W}}(P)$ of $\widetilde{\mathbf{W}}(\Phi)$ in the cover. Consider the functor
$\operatorname{Fuk}_{*}\colon\operatorname{UI}(\Phi)\to^{**}\operatorname{DG}$
which sends each Weinstein sector to the category of modules over the wrapped Fukaya category. The images of morphisms are induced by the pushforward functors from [GPS1] along the inclusions of Weinstein sectors. Due to the covariant functoriality and sectorial descent of wrapped Fukaya categories, $\operatorname{Fuk}_{*}$ gives a cosheaf of categories over $\operatorname{UI}(\Phi)$. Note that $\operatorname{Fuk}_{*}$ preserves compact objects as it is defined before taking module categories. Since $\bar{\pi}$ restricts to $\pi$ and $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a lift of an open cover of $\widetilde{\mathbb{L}}(\Phi)$, the pushforward
@@
There is an isomorphism
$\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\bar{\pi}_{*}\operatorname{Fuk}_{*}.$
@@
of $\operatorname{Coh}_{!}$ with $\mathbf{T}$ from [GS23, Section 3] gives a constructible cosheaves of categories over $\Phi$. Recall that $\mathbf{T}(S^{(k)})$ is the toric variety associated with the fan $\Sigma_{S^{(k)}}\subset M_{S^{(k)},\mathbb{R}}$, where $(\Sigma_{S^{(k)}},M_{S^{(k)}})$ is the image of $S^{(k)}$ under the functor $\operatorname{Exit}(\Phi)^{op}\to(\operatorname{Fan}^{\twoheadrightarrow})^{op}$ equipped with $\Phi$. Recall also that the values of $\operatorname{Coh}_{!}\circ\mathbf{T}$ are the module categories over the dg categories of coherent sheaves on toric varieties and the images of morphisms are the pushforwards along closed immersions.
Lemma 4.14 ([GS23, Proposition 3.10]).
The colimit
$\mathbf{T}(\Phi)=\operatorname{colim}_{\operatorname{Exit}(\Phi)^{op}}\mathbf{T}(S^{(k)})$
exists as an algebraic space.
Lemma 4.15 ([GS23, Proposition 3.14]).
The canonical functor
@@
is an equivalence.
Now, it is straightforward to describe HMS in terms of cosheaves of categories over $\Phi$.
Lemma 4.16 (cf. [GS23, Theorem 5.3]).
There is an isomorphism
$\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}.$
@@
The claim follows immediately from [GS23, Theorem 5.3]. ∎
Theorem 4.17 (cf. [GS23, Corollary 5.8]).
There is an isomorphism
$\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}.$
Proof.
Concatenate Lemma 4.13 and Lemma 4.16. ∎
Taking the global section, we obtain
Corollary 4.18 (cf. [GS23, Theorem 5.4]).
There is an equivalence
@@ 5.1. Mirror pair for very affine hypersurfaces
To a very affine hypersuface $H\subset M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}\cong(\mathbb{C}^{*})^{d+1}$ satisfying certain assumptions on the defining Laurent polynomial, the toric boundary divisor $\partial\mathbf{T}_{\Sigma}$ gives its mirror [GS22]. Here, $\mathbf{T}_{\Sigma}$ is the associated toric stack with a smooth quasiprojective stacky fan $\Sigma\subset M_{\mathbb{R}}\cong\mathbb{R}^{d+1}$. Recall that $\Sigma$ is determined by the defining polynomial of $H$. Under the assumptions, its stacky primitives span a convex lattice polytope $\Delta^{\vee}$ containing the origin and $\Sigma$ defines an adapted star-shaped triangulation $\mathcal{T}$ of $\Delta^{\vee}$. Below, we review the proof of the following statement.
Theorem 5.1 ([GS22, Theorem 1.0.1]).
Assume that the Newton polytope $\Delta^{\vee}$ contains the origin and its triangulation $\mathcal{T}$ is star-shaped. Then there is an equivalence
@@
Due to the assumptions, one can easily construct a global skeleton $\operatorname{Core}(H)$ of $H$. Recall that we restrict to $H$ the canonical Weinstein structure on $M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}$. In the sequel, for simplicity we will assume that $\Sigma$ is an ordinary simplicial fan. When $\Sigma$ is a smooth quasiprojective stacky fan, we will replace the $A$-model with its finite cover and the $B$-model with the associated toric stack. The skeleton $\operatorname{Core}(H)$ can be obtained by gluing skeleta of tailored pants $\tilde{P}_{d}$ computed by Nadler [Nad], after transported to the ones in the pants decomposition of $H$. As explained in [Nad], we may identify the skeleton $\operatorname{Core}(\tilde{P}_{d})$ with the ideal boundary of the FLTZ skeleton associated with the toric variety $\mathbb{A}^{d+1}$. Hence we may identify $\operatorname{Core}(H)$ with $\partial_{\infty}\mathbb{L}(\Sigma)$ for the FLTZ skeleton $\mathbb{L}(\Sigma)$ associated with $\mathbf{T}_{\Sigma}$.
Lemma 5.2 ([GS22, Corollary 4.3.2]).
Assume that the Newton polytope $\Delta^{\vee}$ contains the origin and its triangulation $\mathcal{T}$ is star-shaped. Then $\partial_{\infty}\mathbb{L}(\Sigma)$ has an open cover by the subsets $\mathbb{L}(\Sigma/\sigma)\times\partial_{\infty}\sigma$ anti-indexed by the poset of nonzero cones in $\Sigma$.
Proof.
Let $\sigma\subset\sigma^{\prime}\subset\tau$ be nonzero cones in $\Sigma$. Then $\tau/\langle\sigma\rangle,\tau/\langle\sigma^{\prime}\rangle$ give cones in the normal fans $\Sigma/\sigma,\Sigma/\sigma^{\prime}$ with the isomorphisms
$(\tau/\langle\sigma\rangle)^{\perp}\times\tau/\langle\sigma\rangle\times\sigma\cong(\tau)^{\perp}\times\tau\cong(\tau/\langle\sigma^{\prime}\rangle)^{\perp}\times\tau/\langle\sigma^{\prime}\rangle\times\sigma^{\prime}.$
Note that subtori $(\tau/\langle\sigma\rangle)^{\perp},(\tau/\langle\sigma^{\prime}\rangle)^{\perp}$ of respective tori $\widehat{M}/\langle\sigma\rangle,\widehat{M}/\langle\sigma^{\prime}\rangle$ coincide with $\tau^{\perp}\subset\widehat{M}$. Hence we obtain an open inclusion
$\mathbb{L}(\Sigma/\sigma^{\prime})\times\sigma^{\prime}\hookrightarrow\mathbb{L}(\Sigma/\sigma)\times\sigma.$
Now, the claim follows from $\partial_{\infty}(\mathbb{L}(\Sigma/\sigma)\times\sigma)\cong\mathbb{L}(\Sigma/\sigma)\times\partial_{\infty}\sigma$. ∎
To each $\sigma\in\Sigma$, let $\overline{O(\sigma)}$ be the closure of the corresponding toric orbit $O(\sigma)$. It is the toric variety associated with $\Sigma/\sigma$. Here, we take its reduced induced scheme structure. If $\sigma\subset\tau$, then there is a closed immersion $\overline{O(\tau)}\hookrightarrow\overline{O(\sigma)}$ induced by the quotient $\Sigma/\sigma\to\Sigma/\tau$. Also, we have $\overline{O(\sigma)}\cap\overline{O(\sigma^{\prime})}=\overline{O(\sigma\wedge\sigma^{\prime})}$ where $\sigma\wedge\sigma^{\prime}$ denotes the smallest cone in $\Sigma$ containing both $\sigma$ and $\sigma^{\prime}$ if such a cone exists, otherwise $\emptyset$. Recall that by definition $\partial\mathbf{T}_{\Sigma}$ is the union of the nontrivial toric orbit closures. Scheme-theoretically, this amounts to
Lemma 5.3 ([GS22, Lemma 3.4.1]).
The colimit
$\partial\mathbf{T}_{\Sigma}=\operatorname{colim}_{0\neq\sigma\in\Sigma}\overline{O(\sigma)}$
exists as an algebraic space.
From the above lemma it follows
@@
Now, one can easily obtain (5.1). Indeed, by [GPS3, Theorem 1.4] and Lemma 5.2 we have
@@
As explained in the proof of [GS22, Theorem 7.4.1], we have

### v3

@@ 3.3. Construction
Suppose that $\Phi$ is not closed and, for instance, there is a half-open $1$-stratum II adjacent to $P$. Recall that in this case the handle attachment $\mathbf{W}(\Phi_{1})\#_{\mathcal{L}_{1}}(T^{*}\widehat{M}_{I}\times T^{*}I_{\circ})$ is not precisely a Liouville manifold-with-boundary, as one needs to modify the canonical Liouville form of $T^{*}I_{\circ}$ near $\partial T^{*}I_{\circ}$ making it convex. This modification, which is not necessary if $I$ were closed, yields a stop on the newly formed ideal boundary of the handle attachment.
@@
We adapt the idea in [GPS2, Example 1.34]. Again, assume that $\Phi$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata. Finiteness of the cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ follows from the condition $(i)$. By construction each $\widetilde{\mathbf{W}}(P)$ has no corners and $\partial\widetilde{\mathbf{W}}(\Phi)$ is empty. Moreover, $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ intersect transversely for any distinct pair $P,P^{\prime}$ of 0-strata unless they are disjoint. By Lemma 3.3 each $\widetilde{\mathbf{W}}(P)$ is a Weinstein sector. Hence the collection of the hypersurfaces $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial if there exist functions
@@
Consider the functions $y_{P,I}$ in the proof of Lemma 3.3. Pushing forward $X_{y_{P,I}}$ by $\bar{\pi}$, we obtain a collection of vector fields on neighborhoods of $\bar{\pi}(H_{P,I})$. Here, by a vector field on a subset of $\Phi$ we mean the gluing of vector fields on submanifolds of $\Phi_{k}\setminus\Phi_{k-1}$ for $k=1,\ldots,n$. Take their deformations around the corners to concatenate $\bar{\pi}_{*}(X_{y_{P,I}})$ so that translation of the result gives a vector field $\tilde{X}_{P}$ on a neighborhood $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$, which is outward pointing along $\partial^{2}\widetilde{\mathbf{W}}(P)$. Recall that $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P)),\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P^{\prime}))$ intersect transversely for any distinct pair $P,P^{\prime}$ of 0-strata unless they are disjoint. Since $\Phi$ has only finitely many 0-strata, we may deform further $\tilde{X}_{P}$ simultaneously so that they in addition become tangent to $\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ on the intersections. We use the same symbol $\tilde{X}_{P}$ to denote these deformations. We define $\tilde{y}_{P}$ as the Hamiltonian lifts of $\tilde{X}_{P}$, which make sense due to the construction of $\widetilde{\mathbf{W}}(\Phi)$ and our definition of a vector field on a subset of $\Phi$.
@@
The construction does not work without modification of $\widetilde{\mathbf{W}}(\Phi)$, since in general $\partial\widetilde{\mathbf{W}}(\Phi)$ is nonempty and intersects $\partial^{2}\widetilde{\mathbf{W}}(P)$.
@@ 3.4. Sectorial descent and HMS over fanifolds
which is obtained from $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)}\colon\operatorname{Exit}(\Phi)\to^{*}\operatorname{DG}^{*}$ in [GS23, Section 4.5] by taking adjoints of the images of the morphisms. By definition both $\operatorname{fsh}_{*}$ and $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)*}$ are constructible cosheaves of categories over $\Phi$. From [GS23, Proposition 4.34] we obtain an isomorphism
@@
Third, let $\operatorname{UI}(\Phi)$ be the category of unions of intersections of $\widetilde{\mathbf{W}}(P)$ in the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ from Lemma 3.6. Morphisms are given by inclusions of Weinstein sectors. Introduce to $\operatorname{UI}(\Phi)$ a topology generated by maximal proper Weinstein subsectors $\widetilde{\mathbf{W}}(P)$ of $\widetilde{\mathbf{W}}(\Phi)$ in the cover. Consider the functor $\operatorname{Fuk}_{*}\colon\operatorname{UI}(\Phi)\to^{**}\operatorname{DG}$ which sends each Weinstein sector to the category of modules over the wrapped Fukaya category. The images of morphisms are induced by the pushforward functors from [GPS1] along the inclusions of Weinstein sectors. Due to the covariant functoriality and sectorial descent of wrapped Fukaya categories, $\operatorname{Fuk}_{*}$ gives a cosheaf of categories over $\operatorname{UI}(\Phi)$. Note that $\operatorname{Fuk}_{*}$ preserves compact objects as it is defined before taking module categories. Since $\bar{\pi}$ restricts to $\pi$ and $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a lift of an open cover of $\widetilde{\mathbb{L}}(\Phi)$, the pushforward
@@
There is an isomorphism $\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\bar{\pi}_{*}\operatorname{Fuk}_{*}$.
@@
of $\operatorname{Coh}_{!}$ with $\mathbf{T}$ from [GS23, Section 3] gives a constructible cosheaves of categories over $\Phi$. Recall that $\mathbf{T}(S^{(k)})$ is the toric variety for the fan $\Sigma_{S^{(k)}}\subset M_{S^{(k)},\mathbb{R}}$, where $(\Sigma_{S^{(k)}},M_{S^{(k)}})$ is the image of $S^{(k)}$ under the functor $\operatorname{Exit}(\Phi)^{op}\to(\operatorname{Fan}^{\twoheadrightarrow})^{op}$ equipped with $\Phi$. Recall also that the values of $\operatorname{Coh}_{!}\circ\mathbf{T}$ are the module categories over the dg categories of coherent sheaves on toric varieties and the images of morphisms are the pushforwards along closed immersions. By [GS23, Proposition 3.10] the colimit $\mathbf{T}(\Phi)=\operatorname{colim}_{\operatorname{Exit}(\Phi)^{op}}\mathbf{T}(S^{(k)})$ exists as an algebraic space. Moreover, the canonical functor
@@
is an equivalence in [GS23, Proposition 3.14].
Theorem 3.9 (cf. [GS23, Corollary 5.8]).
There is an isomorphism $\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$.
@@
From [GS23, Theorem 5.3] we obtain an isomorphism $\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$. Combined with Lemma 3.8, it yields the desired isomorphism. ∎
Taking the global section, we recover the equivalence [GS23, Theorem 5.4]
@@ 4.1. Mirror pair for very affine hypersurfaces
Let $H\subset M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}\cong(\mathbb{C}^{*})^{d+1}$ be a very affine hypersurface. Assume that, via the standard correspondence, its defining Laurent polynomial determines a smooth quasiprojective stacky fan $\Sigma\subset M_{\mathbb{R}}\cong\mathbb{R}^{d+1}$. Assume further that its stacky primitives span a convex lattice polytope $\Delta^{\vee}$ containing the origin and $\Sigma$ defines an adapted star-shaped triangulation $\mathcal{T}$ of $\Delta^{\vee}$. Let $\mathbf{T}_{\Sigma}$ be the toric stack for $\Sigma$. By [GS22, Theorem 1.1] for its toric boundary divisor $\partial\mathbf{T}_{\Sigma}$ there is an equivalence
@@
Below, we briefly review the proof. Due to the assumptions, one can construct a global skeleton $\operatorname{Core}(H)$ of $H$. Recall that we restrict to $H$ the canonical Weinstein structure on $M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}$. In the sequel, for simplicity we will assume that $\Sigma$ is an ordinary simplicial fan. When $\Sigma$ is a smooth quasiprojective stacky fan, we will replace the $A$-model with its finite cover and the $B$-model with the associated toric stack. The skeleton $\operatorname{Core}(H)$ can be obtained by gluing skeleta of tailored pants $\tilde{P}_{d}$ computed by Nadler [Nad], after transported to the ones in the pants decomposition of $H$. As explained in [Nad], we may identify the skeleton $\operatorname{Core}(\tilde{P}_{d})$ with the ideal boundary of the FLTZ skeleton associated with the toric variety $\mathbb{A}^{d+1}$. Hence we may identify $\operatorname{Core}(H)$ with $\partial_{\infty}\mathbb{L}(\Sigma)$ for the FLTZ skeleton $\mathbb{L}(\Sigma)$ associated with $\mathbf{T}_{\Sigma}$. Moreover, by [GS22, Corollary 4.8] the skeleton $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Sigma)$ has an open cover by the subsets $\mathbb{L}(\Sigma/\sigma)\times\partial_{\infty}\sigma$ anti-indexed by the poset of nonzero cones in $\Sigma$.
To each $\sigma\in\Sigma$, let $\overline{O(\sigma)}$ be the closure of the corresponding toric orbit $O(\sigma)$. It is the toric variety for $\Sigma/\sigma$. Here, we take its reduced induced scheme structure. If $\sigma\subset\tau$, then there is a closed immersion $\overline{O(\tau)}\hookrightarrow\overline{O(\sigma)}$ induced by the quotient $\Sigma/\sigma\to\Sigma/\tau$. Also, we have $\overline{O(\sigma)}\cap\overline{O(\sigma^{\prime})}=\overline{O(\sigma\wedge\sigma^{\prime})}$ where $\sigma\wedge\sigma^{\prime}$ denotes the smallest cone in $\Sigma$ containing both $\sigma$ and $\sigma^{\prime}$ if such a cone exists, otherwise $\emptyset$. By definition $\partial\mathbf{T}_{\Sigma}$ is the union of the nontrivial toric orbit closures. Scheme-theoretically, this amounts to the existence of the colimit
(4.2) $\partial\mathbf{T}_{\Sigma}=\operatorname{colim}_{0\neq\sigma\in\Sigma}\overline{O(\sigma)}$
as an algebraic space [GS22, Lemma 3.5]. Then we have
@@
By [GPS3, Theorem 1.4] and [GS22, Corollary 4.8] we have
@@
As explained in the proof of [GS22, Theorem 7.13], we have